OpenStudy (anonymous):

Find the solutions to this equation algebraically using the Fundamental Theorem of Algebra, the Rational Root Theorem, Descartes' Rule of Signs, and the Factor Theorem to 204.75=x^(2)-8x+13.75

OpenStudy (anonymous):

Are you not allowed to use the quadratic formula?

OpenStudy (anonymous):

i dont know. i was just told to do these formulas. i am sure my teacher wouldnt mind if i did the quadratic formula though. even though i dont understand that either.

OpenStudy (anonymous):

Those theorems that you stated don't really help you solve the problem. The answer is:\[4\pm3\sqrt{23}\]

OpenStudy (anonymous):

ok thank you would u be able to show me the steps?

OpenStudy (anonymous):

Do you want a more detailed explanation?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

\[x ^{2}-8x+13.75=204.75\]\[x ^{2}-8x-191=0\]\[\frac{ 8\pm \sqrt{64-4(-191)} }{ 2 }\]\[\frac{ 8\pm \sqrt{828} }{ 2 }\]\[\frac{ 8\pm6\sqrt{23} }{ 2 }\]\[4\pm3\sqrt{23}\]

OpenStudy (anonymous):

The third step I put everything into the quadratic formula

OpenStudy (anonymous):

oo ok thank you..i think i understand it better now.

OpenStudy (anonymous):

your wlecome